OpenStudy (anonymous):

The sum of two numbers is 5 and their difference is 2. Which of the following could be the difference of their squares

hartnn (hartnn):

x+y = 5 x-y=2 just multiply them, what u get ?

hartnn (hartnn):

do you know why we are 'multiplying' here ?

OpenStudy (anonymous):

Because of the squares?

hartnn (hartnn):

difference of squares = \(x^2-y^2=(x+y)(x-y)\) thats why :)

OpenStudy (anonymous):

So multiply 5 and 2?

hartnn (hartnn):

yup.

hartnn (hartnn):

don't tell me 10 isn't in choices... :P

OpenStudy (anonymous):

No, you're right! :D

hartnn (hartnn):

your welcome ^_^